OpenStudy (anonymous):

A math test is made up of 5-point questions and 10-point questions. There are 24 questions on the test. The total number of possible points is 150 points. A. Write a system of linear equations to represent the problem. Use f for 5-point questions and t for 10-point questions. B. Find the intercepts of the graph of each equation you wrote in part a. State your answers in complete sentences. C. Use the intercepts from part b. to graph the system of linear equations. Label each axis to show what it represents. Label each line. What is the total value of all the 10-point questions on t

OpenStudy (anonymous):

a, F + T = 24 and 5F + 10T = 150

OpenStudy (anonymous):

do you see that ?

OpenStudy (anonymous):

A. F+T=24 and, 5F+ 10T=150 B. No idea, lol.

OpenStudy (anonymous):

The intercepts of a single equation are where it crosses the x and y axis. Thus set x to zero and y to zero to get those.

OpenStudy (bahrom7893):

^what pjs said: B) 1) F+T=24 uhmm which one do we pick for y? I'm thinking t should be x

OpenStudy (bahrom7893):

anyway t is x, f is y: y+x=24 y = -x+24 intercepts are: (24; 0) and (0; 24)

OpenStudy (bahrom7893):

5F + 10T = 150 5y + 10x =150 y +2x = 30 y = -2x+30 Intercepts: (15;0); (0; 30)